OpenStudy (kidthatbro8):

Simplify. square root of 7 over fourth root of 7 7 to the power of negative 3 over 4 7 to the power of 1 over 4 7 to the power of 3 over 4 7

OpenStudy (anonymous):

\[\sqrt[b]{x^a} = x^{\frac{ a }{ b }}\] \[\frac{ x^a }{ x^b } = x^{a-b}\]

OpenStudy (anonymous):

Use these rules.

OpenStudy (kidthatbro8):

i'm not sure how to get the problem to look like that, though.

OpenStudy (anonymous):

Let's break it down then, what is the square root of 7 according to the first rule?

OpenStudy (jdoe0001):

\(\Large { \cfrac{\sqrt[2]{7}}{\sqrt[4]{7}} \\ \quad \\ \quad \\ a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}} \\\quad \\ % ----------------------------- \cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}}= \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}}\implies a^{-\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ \cfrac{\sqrt[2]{7^1}}{\sqrt[4]{7^1}}\implies \cfrac{7^{\frac{1}{2}}}{7^{\frac{1}{4}}}\implies \cfrac{7^{\frac{1}{2}}}{1}\cdot \cfrac{1}{7^{\frac{1}{4}}}\implies 7^{\frac{1}{2}}\cdot 7^{-\frac{1}{4}} }\)

OpenStudy (anonymous):

Or just look at this very descriptive explanation.

OpenStudy (jdoe0001):

so... simpify that... recall your exponent rules

OpenStudy (kidthatbro8):

it is a very good explanation. thank you. how would i solve it ?

OpenStudy (jdoe0001):

use the exponent rules, check your material on that or just check this rules set http://ts3.mm.bing.net/th?id=HN.607986448326853914&pid=15.1&P=0

OpenStudy (jdoe0001):

when multiplying two bases, different exponent keep the base, sum the exponents

OpenStudy (jdoe0001):

well when multiplying two SAME bases, different exponent keep the base, sum the exponents

OpenStudy (kidthatbro8):

so...i would be adding the exponents, then ? and i'd need to make the denominators the same ?

OpenStudy (jdoe0001):

yes... since they're fractions so you'd be adding the fractions, yes, so yo'd get the LCD \(\large { \cfrac{\sqrt[2]{7^1}}{\sqrt[4]{7^1}}\implies \cfrac{7^{\frac{1}{2}}}{7^{\frac{1}{4}}}\implies \cfrac{7^{\frac{1}{2}}}{1}\cdot \cfrac{1}{7^{\frac{1}{4}}}\implies 7^{\frac{1}{2}}\cdot 7^{-\frac{1}{4}}\implies 7^{\frac{1}{2}-\frac{1}{4}} }\)

OpenStudy (kidthatbro8):

the lcd is 8, right ?

OpenStudy (jdoe0001):

you could use 8, sure, is not the LCD, but doesn't matter since it'd be simplified anyway

OpenStudy (kidthatbro8):

oh, okay, so then the answer is 7^1/4, right ?

OpenStudy (jdoe0001):

yeap